How quickly can you do this? Fill appropriate sign. ( ‘’) 
6 / 10 ☐ 4 / 5

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

6 / 10 = (3 × 2) / (5 × 2)
= 3 / 5
Between 3 / 5 and 4 / 5
Both have same denominators. So, the fraction having greater numerator will be greater
∴6 / 10 < 4 / 5

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We have fractions with unequal denominator,
In order to compare, we'll need to make denominator equal.
6/10 can be written as 3/5
 
Now, the denominator is equal.
So, 4/5 > 3/5
 
Hope it helps!
Thank you